The larger of two numbers is 5 more than the smaller.If their product is 84,what are the numbers?
Answer by Cromlix(4381)

Smaller number = x
Larger number = x + 5
x(x + 5) = 84
x^2 + 5x - 84 = 0
(x - 7)(x + 12) = 0
x - 7 = 0
x = 7
x + 12 = 0
x = -12 (No answer as x < 0)
Smaller number = 7
Larger number = 12
Hope this helps
:-)